A compound is found to contain 50.05 % sulfur and 49.95 % oxygen by mass.

What is the empirical formula for this compound?

A compound is found to contain 6.360 % silicon, 36.18 % bromine, and 57.46 % iodine by mass.

What is the empirical formula for this compound?

A compound is found to contain 7.523 % phosphorus and 92.48 % iodine by mass.

What is the empirical formula for this compound?
Answer #1

1)

we have mass of each elements as:

S: 50.05 g

O: 49.95 g

Divide by molar mass to get number of moles of each:

S: 50.05/32.07 = 1.561

O: 49.95/16.0 = 3.122

Divide by smallest to get simplest whole number ratio:

S: 1.561/1.561 = 1

O: 3.122/1.561 = 2

So empirical formula is:SO2

Answer: SO2

2)

we have mass of each elements as:

Si: 6.36 g

Br: 36.18 g

I: 57.46 g

Divide by molar mass to get number of moles of each:

Si: 6.36/28.09 = 0.2264

Br: 36.18/79.9 = 0.4528

I: 57.46/126.9 = 0.4528

Divide by smallest to get simplest whole number ratio:

Si: 0.2264/0.2264 = 1

Br: 0.4528/0.2264 = 2

I: 0.4528/0.2264 = 2

  

So empirical formula is:SiBr2I2

Answer: SiBr2I2

3)

we have mass of each elements as:

P: 7.523 g

I: 92.48 g

Divide by molar mass to get number of moles of each:

P: 7.523/30.97 = 0.2429

I: 92.48/126.9 = 0.7288

Divide by smallest to get simplest whole number ratio:

P: 0.2429/0.2429 = 1

I: 0.7288/0.2429 = 3

So empirical formula is:PI3

Answer: PI3